OpenStudy (anonymous):

Solve the equation 15-x = 24. Round your answer to the nearest hundredth. A. x ≈ 2.45 B. x ≈ -1.29 C. x ≈ -1.17 D. x ≈ 1.67

OpenStudy (shamim):

\[15-x=24\]

OpenStudy (shamim):

\[or, 24=15-x\]

OpenStudy (shamim):

\[or, x=15-24\]

OpenStudy (shamim):

\[or, x=-9\]

OpenStudy (anonymous):

nope, its 15-x=24